# TopoCode: Topologically Informed Error Detection and Correction in Communication Systems

### v2

@@ III Error Detection and Correction
In order to solve Eq. (4), we can use gradient decent. It is essential to ensure that the filtration and the $p$-Wasserstein distance computation are differentiable. The optimization of persistent homology functions has been studied in [10]. The persistence diagram is formed by BD pairs which are generated by simplicial complexes. The simplicial complex is generated based on distance functions which are differentiable. It is also proved in [10] that the $p$-Wasserstein distance function is Lipschitz and definable in any o-minimal structure. As a result, we can solve Eq. (4) using the following steps. First, ${\hat{\bm{X}}}$ and ${\bm{X}}$ are normalized by dividing their maximum values before obtaining the persistence diagram. Second, given the target persistence diagram $D(f_{l}({\bm{X}}))$, we define the optimization loss function as
$\sum_{i}\min(|{\hat{\bm{X}}}_{[i]}|,|1-{\hat{\bm{X}}}_{[i]}|),$ (5)
IV Simulations and Discussion
@@
Next, we show an example of error correction using MNIST dataset. We compare TopoCode with existing error correction codes including LDPC and Convolutional code. The code rates of LDPC and Convolutional codes are 0.7 and 2/3, respectively. For LDPC and Convolutional codes, the image is converted to binary data, encoded, modulated using BPSK and added with Gaussian noises and the SNR is 3dB. Then, the received data is demodulated, decoded and converted back to decimal numbers. For TopoCode, the persistence diagram is obtained using the image and concatenated with the binarized image. Then, the image is modulated using BPSK and added with Gaussian noises with the same SNR. The received data is demodulated and converted back to decimal numbers. Then, the error correction solution in the previous section is used to recover the image.
@@
In Fig. 7, we show five recovered images. The uncoded image is the one without any coding and the TopoCode is concatenated with uncoded image. The TopoCode image is the recovered one after applying the error correction. As we can see, we can remove the majority of noises and recover the topology of the data by optimizing the persistence diagram. However, since the Cubical Complex that is used to generate persistence diagrams cannot effectively process pixels on the boundary [8], there are remaining noises. In the TopoCode without Boundary image, we removed the boundary pixels and the impact of noises further reduces. The decoded images using LDPC and Convolutional Code are also shown which contain significant noises. This example visually shows that TopoCode is robust in maintaining data topology in presence of noises, while existing bit-level coding schemes cannot effectively capture data topology or correcting errors based on the content. To further validate the proposed solution, we simulate the transmission of 200 images in MNIST and Omniglot datasets and evaluate the recovered images’ PSNR, SSIM, Wasserstein distance, and the code length. The three approaches use the same data and thus the packet length is the length of data and channel coding. In Fig. 8 and Fig. 9, we show the results of TopoCode, TopoCode without boundary, LDPC, and Convolutional code. As we can see from the figure, the TopoCode achieves higher PSNR and SSIM performance in the low-SNR regime. As the SNR increases, the LDPC and Convolutional Code perform well. TopoCode always shows small Wasserstein distance and short code length.
@@
The error correction gain of TopoCode at high SNR regime is not as good as that at low SNR regime. TopoCode can only recover data’s topological information, while non-topological information cannot be captured which limits the performance of TopoCode. Although the MNIST and Omniglot datasets contain simple topological information and the persistence diagrams are small which leads to short code length, there are various algorithms to reduce the size of TopoCode, such as Persistence Landscape, quantification of persistence diagram, and Persistence Image [6].
@@ V Conclusion
In this letter, we propose TopoCode which utilizes data topology to detect and correct errors in received data within communication systems. Error detection is achieved through total persistence, enabling the identification of significant errors while filtering out insignificant ones. For error correction, TopoCode employs persistence function optimization to restore the data topology using persistence diagram. Simulation results demonstrate that TopoCode offers unique advantages in error detection and correction at the message level.

### v3

@@ IV Simulation and Discussion
Next, we show an example of error correction using MNIST dataset. We compare Topocode with existing error correction codes including Low-Density Parity-Check (LDPC) and Convolutional codes. The regular LDPC code uses a codeword length of 1000 and degrees of (5,20). The code rate of Convolutional code is 2/3. For LDPC and Convolutional codes, the image is converted to binary data, encoded, modulated using BPSK and added with Gaussian noises and the SNR is 3dB. Then, the received data is demodulated, decoded and converted back to decimal numbers. For Topocode, the PD is obtained using the image and concatenated with the binarized image. Then, the image is modulated using BPSK and added with Gaussian noises with the same SNR. The received data is demodulated and converted back to decimal numbers. Then, the error correction solution in the previous section is used to recover the image.
@@
In Fig. 7, we show five recovered images. The uncoded image is the one without any coding and the Topocode is concatenated with uncoded image. The Topocode image is the recovered one after applying the error correction. As we can see, we can remove the majority of noises and recover the topology of the data by optimizing the PD. However, since the Cubical Complex that is used to generate PDs cannot effectively process pixels on the boundary [8], there are remaining noises. In the Topocode without Boundary image, we removed the boundary pixels and the impact of noises further reduces. The decoded images using LDPC and Convolutional Code are also shown which contain significant noises. This example visually shows that Topocode is robust in maintaining data topology in presence of noises, while existing bit-level coding schemes cannot effectively capture data topology or correcting errors based on the content. Also, there are 15 BD pairs in this example which corresponds to a Topocode length of 256 bits using 8-bit unsigned integer. Compared with the data length $28\times 28\times 8=6,272$ bits, the code length is only around 4% of source data. To further validate the proposed solution, we simulate the transmission of 200 images in MNIST and Omniglot datasets and evaluate the recovered images’ PSNR, SSIM, Wasserstein distance, and the code length. The three approaches use the same data and thus the packet length is the length of data and channel coding. In Fig. 8 and Fig. 9, we show the results of Topocode, Topocode without boundary, LDPC, and Convolutional code. As we can see from the figure, the Topocode achieves higher PSNR and SSIM performance in the low-SNR regime. As the SNR increases, the LDPC and Convolutional Code perform well. Topocode always shows small Wasserstein distance and short code length.
@@
The error correction gain of Topocode at high SNR regime is not as good as that at low SNR regime. In the high SNR regime, only a small number of bit errors exists in the received data. Existing error detection and correction codes can effectively detect and correct them. However, if these errors do not affect the topological information in the data, Topocode cannot detect or correct them. Moreover, another limitation of Topocode is that it can only recover data’s topological information, while non-topological information cannot be captured. Although the MNIST and Omniglot datasets contain simple topological information and the PDs are small which leads to short code length, there are various algorithms to reduce the size of Topocode, such as Persistence Landscape, quantification of PD, and Persistence Image [6]. This letter only uses images, but Topocode can be extended to other data modalities, including text, time-series data, and point cloud data. Existing TDA solutions can effectively obtain PDs from these modalities and the error detection and correction are the same as the proposed framework in this letter.
@@ V Conclusion
In this letter, we propose Topocode which utilizes data topology to detect and correct errors in received data within communication systems. Error detection is achieved through total persistence, enabling the identification of significant errors while filtering out insignificant ones. For error correction, Topocode employs persistence function optimization to restore the data topology using Persistence Diagram (PD). Simulation results demonstrate that Topocode offers unique advantages in error detection and correction at the message level.